OpenStudy (anonymous):

1) Using the domain {-1, 0, 4}, what is the range for the equation x – 2y = 0? 2) Using the domain {1, 3, 0, 5}, what is the range for the equation 7x – 2y = 9? 3) What is the domain of the function f(x) = 3x+4? Make sure to include any domain restrictions 4) What is the domain of the function g(x) = √x-3 ? Make sure to include any domain restrictions. 5) What is the domain of the function g(x) = √4 - 2x ? Make sure to include any domain restrictions.

OpenStudy (sidsiddhartha):

x-2y=0 domain =all real x=-1,0,4 given calculate y from above eq and put values of x to get range

OpenStudy (sidsiddhartha):

@FrankiesaysRelax did u get it?

OpenStudy (anonymous):

Just domain, what about number 2?

OpenStudy (anonymous):

what about range for number 1? I don't know how to get the range

OpenStudy (sidsiddhartha):

in no 2 domain=1,3,0,5 from the eq calculate y y=(7x-9)/2 now put 1,3,0,5 one by one and u'll get range

OpenStudy (anonymous):

I don't need the domain for any of the questions, just the range. I don't know how to get the range

OpenStudy (anonymous):

I cannot do the math for it, this is my 3rd time doing this question I give up trying.

OpenStudy (sidsiddhartha):

u just need to know The domain is the x value which is given The range is the y value u just have to calculate y gor given x

OpenStudy (anonymous):

Please I've been doing this all day, would you just make this easier on the both of us and tell me. I don't mean to be rude but I've been on this all day, my brain is fried.

OpenStudy (anonymous):

Thank you this will help a lot, and the rest?

OpenStudy (sidsiddhartha):

apply the same process

OpenStudy (sidsiddhartha):

do u know what is meant by [a,b]?

OpenStudy (anonymous):

No I don't know anything that's why I'm on here, I just want to know the answers I'm sick of doing this question over and over again.

OpenStudy (sidsiddhartha):

look first u need to know diffrence between (a,b) and [a,b] to solve the 4th question (a,b)=a<x<b but [a,b]=a<=x<=b now in the 4th prob g(x)=sqrt(x-3) x values greater than or equal to 3 give values where you can find the square root if x<3 then there will be a negetive term in the square root it will produce complex number so x>=3 so domain = [3, ∞)

OpenStudy (anonymous):

So my answer for number 4 is range = 3, ∞ for sure?

OpenStudy (sidsiddhartha):

no it's domain

OpenStudy (anonymous):

sorry domain = 3, ∞

OpenStudy (anonymous):

And #4?

OpenStudy (sidsiddhartha):

it's not 3,infinity that's wrong expression it's [3, ∞) it is number #4

OpenStudy (sidsiddhartha):

for no 3 f(x)=3x+4 No restrictions Domain: In words: all real numbers Interval Notation: (-∞,∞)

OpenStudy (anonymous):

If number 4 is not 3, ∞ then why did you include 3?

OpenStudy (anonymous):

Honestly you'd save us both time if you would please just tell me the answers straightforward, not to sound rude.

OpenStudy (sidsiddhartha):

it can be written as all real x; x ≥ 3, or (3 ≤ x < ∞)

OpenStudy (anonymous):

This is number 3?

OpenStudy (sidsiddhartha):

no it's #4 #3 i've mentioned above

OpenStudy (anonymous):

This is only confusing me more! Okay #1: (-1/2, 0, 2) and #3 is =-∞,∞ and #4 is 3 ≤ x < ∞

OpenStudy (anonymous):

Are these right, and what about the others?

OpenStudy (sidsiddhartha):

i have already told u the process now u can do it :)

OpenStudy (anonymous):

Are #1, #3, #4 correct? Is there anybody else that will help me with the last two?

OpenStudy (anonymous):

#1: (-1/2, 0, 2) and #3 is =-∞,∞ and #4 is 3 ≤ x < ∞ Are these correct, can I write these down and not fail?

OpenStudy (sidsiddhartha):

have'nt u understood anything??

OpenStudy (anonymous):

No! I haven't, I've sat here for an hour for five small questions!

OpenStudy (sidsiddhartha):

#2=(-1,6,-9/2,13) #5= 2 ≤ x < ∞

OpenStudy (anonymous):

Okay great thank you, finally. Are there any restriction in any of them